evaluate; 2^square root of x + 3square root of x = 9
----------
This equation can be solved by graphing.
2^(sqrt x) =- 3sqrt(x) = 9

Answer by stanbon(48569)

This equation can be solved by graphing.
2^(sqrt x) = -3sqrt(x) + 9
Using a graphing calculator and letting
Y1= 2^(sqrt x)
Y2= 9-3sqrt (x)
The intersection of these two curves is:
x=3.3185408
Cheers,
Stan H.